# Equal Contribution   Equal Advising https://sirui-xu.github.io/InterDreamer/

### v1

@@ 4.1 Experimental Setup
LLM (##\## of parameters) Q1 Acc $\uparrow$ Q1 Acc $\uparrow$ Q2 Acc $\uparrow$ Q2 Acc $\uparrow$
@@
Baselines. As we are introducing a new task, there is no established baseline available at the current stage. Note that it is unfair to compare our work with concurrent supervised learning approaches [50, 71, 21], and their code is not publicly available. To facilitate our comparisons, we develop various baselines to evaluate both our overall pipeline and its individual components. In the context of high-level planning, we utilize GPT-4 [67] and Llama-2 [95], illustrating the effectiveness of our prompts across different language models. For low-level motion generation control, our baselines include MDM [93], MotionDiffuse [130], ReMoDiffuse [131], and MotionGPT [35], which span a range of text-to-motion approaches trained on HumanML3D [25] and show the generalizability of our framework. Our dynamics model baselines are varied, comprising: InterDreamer with InterDiff [116], which adopts their Interaction Correction module as the dynamics model; InterDreamer without control, which operates object dynamics independently of human motion; and InterDreamer with raw control, utilizing unprocessed human motion to guide the dynamics.
4.2 Quantitative Results
Table 1 presents a comparative analysis of our approach, InterDreamer adopting various text-to-motion models against four counterparts on the BEHAVE [7] dataset. Our approach generalizes across various models and consistently outperforms baselines. Specifically, InterDreamer exhibits superior motion quality, reflected by a significantly lower FID, higher R-Precision, and better diversity, highlighting the benefits of incorporating our planning to reduce the distribution gap in the zero-shot setting.
Figure 7: Results from the interaction retrieval. We demonstrate that our proposed retrieval approach based on handcraft rules can extract diverse and realistic interactions. Of these, one interaction is sampled based on a pre-defined metric for subsequent steps.
Figure 8: Ablation study on the dynamics modeling. Given the text description of “A person walks clockwise while holding a small box with left hand,” our (b) vertex-based control can synthesize consistent contacts, which (a) the baseline fails to do.
In Table 2, comparing our whole pipeline to baselines without explicit control, InterDreamer achieves better interaction quality in terms of CMD and penetration scores, showing the importance of human influence on object motion. Against methods that utilize direct raw human motion or markers [116] for control, our full method demonstrates enhanced performance by offering more fine-grained guidance and extracting generalizable features.
4.3 Qualitative Results
Fig. 3 displays several results guided by the text that we annotate on the BEHAVE dataset [7]. Our method exhibits proficiency in interpreting the textual input and synthesizing dynamic, realistic interactions, despite the absence of training with text-interaction paired data. More importantly, our method can process free-form language descriptions of motion that deviate from our annotation, exhibiting zero-shot generalizability in its performance, as illustrated in Fig. 4, where we selectively use more complex sequences of interactive descriptions that are beyond the scope of the original dataset. Fig. 5 further exemplifies the zero-shot ability of our method that is able to generalize effectively to the CHAIRS dataset [36], despite our dynamics model not being trained on it. Fig. 7 depicts the retrieval procedure, resulting in a diverse set of interactions that are both high-quality and semantically aligned. More experimental results and the user study are presented in Sec. D of supplementary.
@@ 4.4 Ablation Study
Adaptability of high-level planning. Is our framework adaptable across different large language models (LLMs)? As illustrated in Table 3, our analysis contains two types of language models: GPT-4 [67], which is accessible through APIs and operates as a black box model; and Llama-2 [95], which is an open-source model. We demonstrate that language models with large parameters exhibit very high accuracy in responding to questions tailored to our prompts, thereby validating the framework’s adaptability.
Effectiveness of text-to-motion with high-level planning. In consistency with Table 1, Fig. 6 offers a qualitative comparison of text-to-motion results, contrasting outputs with and without augmentation by LLM-revised text descriptions. The comparison shows that motions generated without LLM-enhanced descriptions often fail to correspond with the intended text, if the text with the object description is not in the distribution of training data on HumanML3D [25]. This underscores the LLM’s critical role in bridging the gap in a zero-shot way. On the right-hand side of Fig. 6, we see that planning not only reduces the distributional gap in motion, but also directly reduces it in text. We visualize the CLIP [78] features of descriptions on HumanML3D, our raw annotations, and the annotations processed by high-level planning. The text processed by the planning shows more similarity to the in-distributional text, where the average cosine similarity is $0.932$ over $0.913$ from the raw annotation.
Effectiveness of world model. In the quantitative evaluation, we show that the performance of our pipeline is enhanced by the tailored design of our world model. Table 2 provides additional evidence of this effectiveness by integrating the proposed world model, as interaction correction within the InterDiff framework [116] in the interaction prediction task. This implementation demonstrates enhanced conditionality in the object dynamics modeling across various tasks, attributed to the vertex-level control. Doing so effectively removes the whole-body complexity, most of which tends to be irrelevant to the interaction. Fig. 8 further indicates that our vertex-based control is able to establish consistent interactions over time, while guidance from motion features is not robust.
@@ 5 Conclusion
We introduce the novel task of text-guided 3D human-object interaction generation, and we aim to achieve this without reliance on text-interaction pair data. To this end, we present InterDreamer that decouples interaction dynamics from semantics, where high-level planning and low-level control are introduced to generate semantically aligned human motion and initial object pose, while a world model is responsible for the object dynamics guided by the interaction. Our approach demonstrates promising effectiveness in this novel task, suggesting its considerable potential for various real-world applications.
Limitations. The current utilization of dynamics modeling could be enhanced. A prospective improvement involves incorporating model-based learning techniques, which empower the agent to more effectively interact with the environment and learn a broader range of skills. The generated results may not be physically plausible and hand poses are rough because they are missing from the dataset, but could be improved by integrating a physics simulator.
Supplementary Material
In this supplementary material, we include additional method details and experimental results: (i) We provide a demo video, explained in Sec. A. (ii) We present additional details of interaction retrieval, world model, and optimization in Sec. B. (iii) We provide implementation details and additional information on the experimental setup in Sec. C. (iv) We provide additional qualitative experiments in Sec. D.
@@ A Visualization Video
Beyond the qualitative results presented in the main paper, we include a video on the project website that offers more detailed visualizations of the task, further illustrating the efficacy of our approach. These demos highlight (i) We conduct a qualitative comparison of our approach with existing text-to-HOI work [21, 71] within the framework of supervised learning. Note that as our setting is the zero-shot generation, it is unfair to compare our work with these approaches; we include the comparison here for additional reference. We evaluate our method by directly testing our trained model on the annotated data available from their websites, specifically retrieving their generated videos for direct comparison. Remarkably, even without training on these datasets, our method generates results that demonstrate high-quality interactions. (ii) The smoothness of the interaction sequences generated by our method. (iii) Remarkably, our method is proficient in creating sequences that maintain consistent contact where contact remains largely unchanged during interactions, such as carrying a box and sitting in a chair. (iv) It is even capable of synthesizing complex interactions involving dynamically-changing contact, such as the handover and throwing of objects. (v) We contrast our framework, which incorporates vertex control, against generation processes that rely on raw motion control. The results demonstrate strong generalizability of our method.
@@ B.1 Low-Level Control
Handcraft Interaction Retrieval. In Sec. 3.2 of the main paper, we detail the construction of the interaction database and emphasize the use of body parts and object categories as keys to fetch semantically-aligned contact maps. Same as the main paper, we define a contact map as a list of $K$ index pairs of vertices $\{(d_{h}^{i},d_{o}^{i})\}_{i=1}^{K}$. This section delves into the methodology for outlining an optimization process to generate the object initial pose $\boldsymbol{s}_{1}$ given contact maps and the initial human pose $\boldsymbol{a}_{1}$, and choose one pose based on a predefined metric.
Let $\boldsymbol{v}_{\boldsymbol{h}_{1}}[d_{o}]$ denote the vertex on the surface of the object, and $\boldsymbol{v}_{\boldsymbol{o}_{1}}[d_{h}]$ represent the corresponding vertex on the human body surface, where $d_{o}$ and $d_{h}$ are the indices of vertices. Specifically, to optimize $\boldsymbol{s}_{1}$, the overall optimization objective is given by,
@@
Fitting Loss. To project a contact map to a object pose, we minimize the L2 distance between the human vertices and the object vertices indicated by the contact map,
$E_{\mathrm{fit}}=\sum_{i=1}^{K}\|\boldsymbol{v}_{\boldsymbol{o}_{1}}[d_{o}^{i}]-\boldsymbol{v}_{\boldsymbol{h}_{1}}[d_{h}^{i}]\|_{2}.$ (10)
Contact Loss. We leverage a contact loss to encourage body part to contact the object surface,
$E_{\mathrm{cont}}=\sum_{d_{h}}\min_{d_{o}}\|\boldsymbol{v}_{\boldsymbol{o}_{1}}[d_{o}]-\boldsymbol{v}_{\boldsymbol{h}_{1}}[d_{h}]\|_{2}.$ (11)
where $d_{h}$ here indicates the vertices index in the certain body part.
Penetration Loss. Given the signed-distance field of the human pose $\textbf{sdf}_{\boldsymbol{h}_{1}}$, we employ a penetration loss to penalize the body-object interpenetration,
$E_{\mathrm{pene}}=-\sum_{d_{o}}\min(\textbf{sdf}_{\boldsymbol{h}_{1}}(\boldsymbol{v}_{\boldsymbol{o}_{1}}[d_{o}]),0).$ (12)
The metric for determining the final pose selection is given by the expression $\mathds{1}(E_{\mathrm{pene}}=0)/E_{\mathrm{cont}}$. We sample one pose from the set of poses generated by all contact maps, based on the metric.
Learning-based Interaction Retrieval. Our interaction retrieval can also be achieved by integrating knowledge from several learning-based algorithms. Although being more complicated, the retrieval can be done without handcraft rules. Our pipeline can be divided into followings. (i) Given the text prompt $\boldsymbol{t}$ and the initial human pose $\boldsymbol{a}_{1}$, we synthesize corresponding images via Stable Diffusion[81]. (ii) We follow[28] filter out images with low quality in interaction (iii) An off-the-shelf model LEMON[120] is employed to obtain object affordance and human contact, given the generated image paired with human pose $\boldsymbol{a}_{1}$ and object template. The output $\{(l_{h}^{i})_{i=1}^{M},(l_{o}^{i})_{i=1}^{N}\}$ indicates the contact vert indexes of human and object respectively, and the output $\boldsymbol{T}_{1}$ indicates the estimated object translation, which is used for initialization in the optimization. (iv) To acquire the object pose, we utilize the optimization to minimize the Chamfer distance between the human vertices and the object vertices, indicated by the contact vertices obtained in the last step.
$E_{\mathrm{fit}}=\sum_{j}\min_{k}\|\boldsymbol{v}_{\boldsymbol{o}_{1}}[l_{o}^{k}]-\boldsymbol{v}_{\boldsymbol{h}_{1}}[l_{h}^{j}]\|_{2}.$ (13)
@@ B.2 World Model
In Sec. 3.3 of the main paper, we introduce the input representation for the dynamics model. The architecture of the world is also introduced in detail. In this section, we introduce detail with particular emphasis on scenarios in which the initial state $\boldsymbol{s}_{1}$ is presented.
In the particular instance where the time step $t=1$, the state vector $\boldsymbol{s}_{1}$ encapsulates a single frame. Consequently, we employ two distinct models for dynamics prediction. For predictions originating from the initial state, the history motion encompasses a single time step $H=1$. In contrast, for predictions for subsequent states, the historical interval covering $m$ time steps, where $m$ denoting the frame count per segment.

### v2

@@ 4.3 Qualitative Results
LLM (#\# of parameters) Q1 Acc $\uparrow$ Q1 Acc $\uparrow$ Q2 Acc $\uparrow$ Q2 Acc $\uparrow$
@@
Table 5: Quantitative comparison of text similarity. The text processed by high-level planning is more similar to text in HumanML3D [34] on average, while addressing the distributional gap significantly for challenging out-of-distributional descriptions, compared to text without planning.
[…]
w/o planning $0.913$ $0.838$
w/ planning $\mathbf{0.932}$ $\mathbf{0.927}$
@@ 4.4 Ablation Study
Adaptability of High-Level Planning. Is our framework adaptable across different large language models (LLMs)? As illustrated in Table 4, our analysis contains two types of language models: GPT-4 [87], which is accessible through APIs and operates as a black box model; and Llama-2 [119], an open-source model. We demonstrate that language models with large parameters exhibit very high accuracy in responding to questions tailored to our prompts, validating the framework’s adaptability.
Effectiveness of High-Level Planning with Low-Level Control. In consistency with Table 2, Figure 7 offers a qualitative comparison of text-to-motion results, contrasting results with and without LLM-revised text descriptions. The comparison shows that motion generated without LLM-enhanced descriptions often fails to correspond to the intended text, if the text is too challenging, e.g., not in the distribution of HumanML3D [34], which is used to train text-to-motion models. This underscores the LLM’s critical role in bridging the distribution gap. In Figure 7(b), we visualize the CLIP [97] features of descriptions from HumanML3D, our raw annotations, and those processed by high-level planning. Quantitative evidence is provided in Table 5. Text processed through high-level planning demonstrates greater similarity to the HumanML3D dataset. Additionally, we test on more challenging out-of-distribution text, selecting examples with an average cosine similarity to HumanML3D text of less than $0.85$. High-level planning successfully rephrases these texts, significantly increasing their similarity. For example, in Figure 7(a), the text “Someone can be seen sitting on a yoga ball” has a cosine similarity of $0.874$ to the closest in-distribution text, while the rephrased text after planning, “A person is seated on an object,” achieves a similarity of $0.958$.
Effectiveness of World Model. In the quantitative evaluation, we show that the performance of our framework is enhanced by the tailored design of our world model. Table 1 provides additional evidence of the effectiveness by integrating the proposed world model, as interaction correction within the InterDiff framework [148] in the interaction prediction task. This implementation demonstrates enhanced conditionality in the object dynamics modeling across two tasks, attributed to the vertex-level condition as actions. Doing so effectively removes the whole-body complexity, most of which tends to be irrelevant to the interaction. Figure 8 further indicates that our vertex-based condition can establish consistent interactions over time, while the condition by raw motion is not robust.
@@ 5 Conclusion
We tackle the task of text-guided 3D human-object interaction generation, aiming to accomplish this without relying on paired text-interaction data. To this end, we present InterDreamer that decouples interaction dynamics from semantics, formulating the task as retrieval-augmented generation and Markov decision process, where high-level planning and low-level control are introduced to generate semantically aligned human motion and initial object pose, while a world model is responsible for the object dynamics guided by the interaction. Our approach demonstrates effectiveness in this novel task, suggesting its potential for various real-world applications.
Limitations. The current utilization of dynamics modeling could be enhanced. A prospective improvement involves incorporating model-based learning techniques, which empower the agent to more effectively interact and learn a broader range of skills. The results may not be physically plausible and lead to artifacts in some cases, for example, foot skating. Hand poses are rough because they are missing from the dataset, but could be improved by integrating a physics simulator.
Acknowledgments
We thank Qiusi Zhan for supporting the implementation and evaluation of high-level planning, Jiangwei Yu and Xiyan Xu for their efforts in dataset annotation, and Xiang Li for the discussions. This work was supported in part by NSF Grant 2106825, NIFA Award 2020-67021-32799, the IBM-Illinois Discovery Accelerator Institute, the Toyota Research Institute, and the Jump ARCHES endowment through the Health Care Engineering Systems Center at Illinois and the OSF Foundation. This work used computational resources on NCSA Delta and PTI Jetstream2 through allocations CIS220014, CIS230012, CIS230013, and CIS240311 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, and on TACC Frontera through the National Artificial Intelligence Research Resource (NAIRR) Pilot.
@@ A Visualization Video
Beyond the qualitative results presented in the main paper, we include two demo videos that offer more detailed visualizations of the task, further illustrating the efficacy of our approach. These demos highlight (i) We conduct a qualitative comparison of our approach with existing text-to-HOI work [28, 90] within the framework of supervised learning. Note that as our setting contains no text supervision, it is unfair to compare our work with these approaches; we include the comparison here for additional reference. We evaluate our method by directly testing our trained model on the annotated data available from their websites, specifically retrieving their generated videos for direct comparison. Remarkably, even without training on these datasets, our method generates results that demonstrate high-quality interactions. It is even capable of synthesizing complex interactions involving dynamically-changing contact, such as the handover and throwing of objects.
@@ B.1 Low-Level Control
In this section, we provide additional details on the retrieval based on handcraft rules, which is straightforward and does not require training. We also investigate a learning-based method without relying on handcrafted designs.
Handcraft Interaction Retrieval. In Sec. 3.2 of the main paper, we detail the construction of the interaction database and emphasize the use of body parts and object categories as keys to fetch semantically-aligned contact maps. Same as the main paper, we define a contact map as a list of $K$ index pairs of vertices $\{(d_{h}^{i},d_{o}^{i})\}_{i=1}^{K}$. This section delves into the methodology for outlining an optimization process to generate the object initial pose $\bm{s}_{1}$ given contact maps and the initial human pose $\bm{a}_{1}$, and choose one pose based on a predefined metric.
Let $\bm{v}_{\bm{h}_{1}}[d_{o}]$ denote the vertex on the surface of the object, and $\bm{v}_{\bm{o}_{1}}[d_{h}]$ represent the corresponding vertex on the human body surface, where $d_{o}$ and $d_{h}$ are the indices of vertices. Specifically, to optimize $\bm{s}_{1}$, the overall optimization objective is given by,
@@
Fitting Loss. To project a contact map to an object pose, we minimize the L2 distance between the human vertices and the object vertices indicated by the contact map,
$E_{\mathrm{fit}}=\|\bm{v}_{\bm{o}_{1}}[d_{o}]-\bm{v}_{\bm{h}_{1}}[d_{h}]\|_{2}.$ (2)
Contact Loss. We leverage a contact loss to encourage the body part to contact the object surface in addition to the fitting loss,
$E_{\mathrm{cont}}=\sum_{\tilde{d}_{h}\in\mathcal{T}}\min_{\tilde{d}_{o}}\|\bm{v}_{\bm{o}_{1}}[\tilde{d}_{o}]-\bm{v}_{\bm{h}_{1}}[\tilde{d}_{h}]\|_{2},$ (3)
where $\mathcal{T}=\{\tilde{d}_{h}|\min_{\tilde{d}_{o}}\|\bm{v}_{\bm{o}_{1}}[\tilde{d}_{o}]-\bm{v}_{\bm{h}_{1}}[\tilde{d}_{h}]\|_{2}\leq\epsilon\}$ includes the index of the body part that is close to the object vertex $\bm{v}_{\bm{o}_{1}}[\tilde{d}_{o}]$, where $\epsilon$ is a hyperparameter, $\tilde{d}_{h}$ and $\tilde{d}_{o}$ are vertex indices for human and object, respectively, in addition to the contact map $\{(d_{h}^{i},d_{o}^{i})\}_{i=1}^{K}$.
Penetration Loss. Given the signed-distance field of the human pose $\textbf{sdf}_{\bm{h}_{1}}$, we employ a penetration loss to penalize the body-object interpenetration,
$E_{\mathrm{pene}}=-\sum_{d_{o}}\min(\textbf{sdf}_{\bm{h}_{1}}(\bm{v}_{\bm{o}_{1}}[d_{o}]),0).$ (4)
The metric for determining the final pose selection is given by the expression $\mathds{1}(E_{\mathrm{pene}}=0)/E_{\mathrm{cont}}$. We sample a pose from the set generated by all contact maps, with higher metrics corresponding to higher selection probability.
Learning-Based Interaction Retrieval. Our interaction retrieval can also be achieved by integrating knowledge from several learning-based algorithms. Although being more complicated, the retrieval can be done without handcraft rules. Our framework can be divided into following. (i) Given the text prompt $\bm{t}$ and the initial human pose $\bm{a}_{1}$, we synthesize corresponding images via Stable Diffusion [100]. (ii) We follow [37] filter out images with low quality in interaction. (iii) An off-the-shelf model LEMON [156] is to employed to obtain object affordance and human contact, given the generated image paired with human pose $\bm{a}_{1}$ and object template. The output $\{(l_{h}^{i})_{i=1}^{K},(l_{o}^{i})_{i=1}^{K}\}$ indicates the contact vertex indexes of human and object respectively, and the output $\bm{T}_{1}$ indicates the estimated object translation, which is used for initialization in the optimization. (iv) To acquire the object pose, we utilize the optimization to minimize the Chamfer distance between the human vertices and the object vertices, indicated by the contact vertices obtained in the last step.
$E_{\mathrm{fit}}=\sum_{j}\min_{k}\|\bm{v}_{\bm{o}_{1}}[l_{o}^{k}]-\bm{v}_{\bm{h}_{1}}[l_{h}^{j}]\|_{2}.$ (5)
@@ B.2 World Model
World Model for Initial States. In the particular instance where the timestep $t=1$, the state vector $\bm{s}_{1}$ encapsulates a single frame. Consequently, we employ two distinct models for dynamics prediction. For predictions originating from the initial state, the history motion encompasses a single timestep $H=1$. In contrast, for predictions for subsequent states, the historical interval covering $m$ timesteps, where $m$ denotes the frame count per segment.
World Model for Implicit Geometry Encoding. The input to the world model includes the trajectories of the human vertices (represented by red small spheres in the top-right of Figure 2 of the main paper), along with the vertex-to-object surface vectors. By adding the vertex-to-object surface vectors to human vertices, one can easily obtain the object vertices (shown as blue small spheres in the top-right of Figure 2 of the main paper). Though the network of the world model does not receive this information directly, it can learn to combine these features to derive it when needed.
World Model for Novel Objects. The world model employs “contact vertices” as an input, which includes features derived from the object distance field. These features encompass the human vertex-to-object surface distance and the human vertex velocity relative to the nearest object vertex as introduced in Sec. 3.3 of the main paper, inherently including information related to the object’s shape. This encoding is consistently applied to both training objects from the BEHAVE [7] dataset and novel objects from the CHAIRS [47] and OMOMO [66] datasets.
World Model for Non-Contact Objects. The network can process inputs without contact conditions by adopting an approach similar to ControlNet [171]. The network comprises two components: $\mathcal{G}$ that operates without contact vertex conditions, applicable in scenarios where no contact occurs, and $\mathcal{F}$, akin to the control components in ControlNet, which incorporates contact vertex conditions into the object trajectory when contact is present. When there is no contact, only the unconditional network is utilized. The model is aware of past object motion and thus needs to learn how human interaction affects the object’s state. This includes understanding how objects follow contact positions or normals by $\mathcal{F}$, as well as how they move without contact by $\mathcal{G}$. With the no-contact object motion data provided by BEHAVE [7], the world model (more specifically, $\mathcal{G}$) learns to infer whether the object should free-fall based on its previous velocity or remain on the ground based on its height.
B.3 Optimization
We provide detailed formulations of optimization objectives, complementing Sec. 3.4 in the main paper. For efficiency, we perform optimization sparsely only if the loss is above a threshold to improve the efficiency. Specifically, given the reference interaction sequence $\{\bm{h}_{i}\}_{i=1}^{L}$ and $\{\bm{o}_{i}\}_{i=1}^{L}$ of arbitrary length $L$, derived from previous steps, we apply gradient descent to optimize human pose sequence $\{\bm{h}_{i}^{\ast}\}_{i=1}^{L}$ and object pose sequence $\{\bm{o}_{i}^{\ast}\}_{i=1}^{L}$, using the loss function,
$E_{\mathrm{fit}}=\sum_{i=1}^{L}(\|\bm{h}_{i}^{\ast}-\bm{h}_{i}\|_{1}+\|\bm{o}_{i}^{\ast}-\bm{o}_{i}\|_{1}).$ (7)
$E_{\mathrm{vel}}=\sum_{i=1}^{L-1}(\|\bm{h}_{i+1}^{\ast}-\bm{h}_{i}^{\ast}\|_{1}+\|\bm{o}_{i+1}^{\ast}-\bm{o}_{i}^{\ast}\|_{1}).$ (8)
Contact loss. We leverage a contact loss to encourage the body part to contact the object surface, if they are close to each other in the initial interaction,
$E_{\mathrm{cont}}=\sum_{i=1}^{L}\sum_{d_{h}\in\mathcal{T}_{i}}\min_{d_{o}}\|\bm{v}_{\bm{o}_{i}^{\ast}}[d_{o}]-\bm{v}_{\bm{h}_{i}^{\ast}}[d_{h}]\|_{2},$ (9)
where $\bm{v}_{\bm{h}_{i}^{\ast}}[d_{h}]$ denotes the vertex on the human body surface, and $\bm{v}_{\bm{o}_{i}^{\ast}}[d_{o}]$ represents the corresponding vertex on the surface of the object. And $\mathcal{T}_{i}=\{d_{h}|\min_{d_{o}}\|\bm{v}_{\bm{o}_{i}}[d_{o}]-\bm{v}_{\bm{h}_{i}}[d_{h}]\|_{2}\leq\epsilon\}$ includes the index of reference human vertex $\bm{v}_{\bm{h}_{i}}[d_{h}]$ that is close to the reference object vertex $\bm{v}_{\bm{o}_{i}}[d_{o}]$, where $\epsilon$ is a hyperparameter, $d_{h}$ and $d_{o}$ are vertex indices for human and object, respectively.
Penetration Loss. Given the signed-distance field of the human pose $\textbf{sdf}_{\bm{h}_{i}^{\ast}}$, we employ a penetration loss to penalize the body-object interpenetration,
$E_{\mathrm{pene}}=-\sum_{i=1}^{L}\sum_{d_{o}}\min(\textbf{sdf}_{\bm{h}_{i}^{\ast}}(\bm{v}_{\bm{o}_{i}^{\ast}}[d_{o}]),0).$ (10)